Sound Level Workset Name Sdverstone PH105 Spring 2018 Show all math stops and write wnis wi The sound level in decibels is given by Sound Level Exercise ?-10 where / is the intensity in W/mi and A 10 w I. If a sound level is 0 dB, does this mean tut there is no sound? Explain. 2. Does negative dB have meaning? Explain 3. The intensity of a sound from a motorcycle is twice that of a certain car in W/m2. What is the difference in the sound levels in decibels? 4. If you double your distance from a spherical source of sound, by how much do you reduce the sound level in dB? (Assume that /varies as ii)
Answer #1

1) If I is the intensity of sound,

then,

sound intensity level, beta = 10*log(I/Io)

0 = 10*log(I/10^-12)

0 = log(I/10^-12)

10^0 = I/10^-12

1 = I/10^-12

I = 10^-12

so, o dB means Intensity of sound is 10^-12 W/m^2

2) Yes. negative decible means the sound intensity is less than 10^-12 W/m^2 which can not be heard by human year.

3) given I2 = 2*I1

difference between sound intensity levels,
beta2 - beta1 = 10*log(I2/Io) - 10*log*(I1/Io)

= 10*log(I2/I1)

= 10*log(2*I1/I1)

= 10*log(2)

= 3.01 dB

4) if distance is doubled intensity becomes 1/4 times

I2 = I1/4

beta1 - beta2 = 10*log(I1/Io) - 10*log*(I2/Io)

= 10*log(I1/I2)

= 10*log(I1/(I1/4))

= 10*log(4)

= 6.02 dB

so, the sound intensity level reduces by 6.02 dB